OpenStudy (anonymous):

What is the solution to the inequality below? square root over x arrow pointing to he eft with a line under it

OpenStudy (anonymous):



OpenStudy (anonymous):

i need an answer asap

Parth (parthkohli):

Square both sides and you get \(x \le 25 \)

OpenStudy (anonymous):

may want to write \[0\leq x \leq 25\]

Parth (parthkohli):

@satellite73 You were hundred yesterday. How 99 again?

OpenStudy (anonymous):

i guess i am being punished for all the wrong answers i give